OpenStudy (anonymous):

you have a piece of wire of length 10 ft from which you construct a circle and/or a square in such a way that the total enclosed area is maximal. what radius does the circle have and what side length does the square have to maximize the enclosed area?

OpenStudy (campbell_st):

Let the perimeter of the square me x, then the circle has a circumference of 10 - x the. the length of 1 side of the square is x/4 and area = x^2/16 the circle \[10 - x = 2 \pi r\] so \[(10 -x)/2\pi = r\] Area of the circle is \[A = \pi (10 - x)^2 /(2 \pi)^2\] Area of\[A=(10 - x)^2/4\pi\] the circle is then the sum of the areas is \[A = x^2/16 + 1/2\pi \times (100 - 20x + x^2)\] Next find dA/dx \[dA/dx = x/8 + 1/(4\pi) (-20 + 2x)\] then let the derivative = 0 \[x(1/8 + 2/4\pi) - 20/(4\pi) = 0\] solve for x....